OpenStudy (anonymous):

When Spheres R US ships bags of golf balls, the number of balls in each bag much be within 8 balls of 100. Write a compound inequality and an absolute value inequality for an acceptable number of golf balls "b" in each bag.

OpenStudy (anonymous):

\[A) 96 \le b \le 104 ...|b-100|\le 8\]

OpenStudy (anonymous):

\[B) 92 \le b \le 108...|b-8| \le 100\]

OpenStudy (anonymous):

\[C) 92 \le b \le 108....|b-100| \le 8 \]

OpenStudy (anonymous):

\[D) 96 \le b \le 104 .... |b-4| \le 100\]

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

i am not clear with your question,write clearly.

OpenStudy (anonymous):

When Spheres R US ships bags of golf balls, the number of balls in each bag must be within 8 balls of 100. Write a compound inequality and an absolute value inequality for an acceptable number of golf balls "b" in each bag.

OpenStudy (anonymous):

\[\left| x-b \right|\le a,means -a \le ( x-b) \le a\]

OpenStudy (anonymous):

so it's C?

OpenStudy (anonymous):

?????????????????

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

\[\left| b-100 \right|\le8,-8 \le b-100 \le 8,100-8\le b \le 8+100 or 92 \le b \le 108\] so you are correct.,otherwise i have not understood the statement.

OpenStudy (anonymous):

Ok:) and one LAST thing I promise!

OpenStudy (anonymous):

is it 1.9444444......

OpenStudy (anonymous):

thats what it says..

OpenStudy (anonymous):

they are saying no algebra, sorry ,i am leaving now.i have to go.